Another word problem that has stumped me.
You recently started the paperwork to purchase your new home, and were just notified that you can move into the house in 2 weeks. You decide to hire a moving company, but are unsure which company to choose. You search online and are interested in contacting two companies. Heavy lifting and quick move, to discuss their rates. Heavy lifing chrges an $80 fee plus $35 per hour. Quick Move charges $55 per hour with no additonal fees.
A. which mover provides a better deal for 2 hours of work? How did I arrive at my answer. (clueless)
B. Which mover provided a better deal for 15 hour of work? How did I arrive at that answer?
C. For what values H(hours) does quick Move offer the better deal? Express my answer as an Inequality. Explain hou I reach that answer.
I seem to get mixed up on how to use the right formulas. can you explain each step so I can understand how to do this!

Answer by Earlsdon(6294)

Start by assigning a letter to represent the unknown quantity.
Let H = the number of hours.
Now write the two equations for the total cost of of the move by each of the two movers. HL = Heavy lifting and QM = Quick move..
HL = 35H+80
QM = 55H
A) Let H = 2 hours.
HL = 35(2)+80
HL = 70+80
HL = $150
QM = 55(2)
QM = $110
Quick moves is the better choice for 2 hours work.
B) Let H = 15 hours.
HL = 35(15)+80 You can work this out.
QM = 55(15) This one too.
The smallest answer wins.
C) Here you need to find the value of H in the inequality:
QM < HL so...
55H < 35H+80 Solve fo H. Subtract 35H from both sides.
20H < 80 Divide both sides by 20.
H < 4 hours
Quick move is the better deal for moves less than 4 hours.